OpenStudy (credmond):

Find the first partial derivatives of f(x, y) = sin(x - y) at the point (7, 7).

OpenStudy (anonymous):

Using the chain rule gives\[\frac{ \partial f }{ \partial x } = \cos \left( x-y \right)\]\[\frac{ \partial f }{ \partial y } = -\cos \left( x-y \right)\]Evaluate using x=7 and y=7.

OpenStudy (anonymous):

do you get this?

OpenStudy (credmond):

Thank you for setting it up for me! cos(7-7)= cos(0) = 1 -cos(7-7)= -cos(0) = -1 Is that all you have to do?

OpenStudy (anonymous):

yep

OpenStudy (credmond):

Thank you!! I have one more question similar to it! Find the first partial derivatives of f(x,y) = 4x-2y/4x+2y Can you help me set this up?

OpenStudy (credmond):

I forgot to add at the point (1,1)

OpenStudy (anonymous):

well when you do partial, you must take the derivative of what you want and then treat all other vairables as constants. so what do you want to differentitate with respect to first? x or y?

OpenStudy (credmond):

X

OpenStudy (anonymous):

\[f(x,y)=4x-\frac{ 2y }{ 4x }+2y\] \[\frac{ \partial f }{ \partial x } = \frac{ \partial }{ \partial x }(4x-\frac{ 2y }{ 4x }+2y)\]

OpenStudy (anonymous):

can you do this? treat variable y as a constant

OpenStudy (credmond):

Soooo... Is it 4- 2y/4 +2y ? Or do the y's go to 0?

OpenStudy (credmond):

Because they are a constant?

OpenStudy (anonymous):

nothing goes to zero

OpenStudy (anonymous):

y is simply treated as a constant

OpenStudy (anonymous):

so it belongs with the number terms

OpenStudy (anonymous):

how about if i simplified the function \[f(x,y)=4x-\frac{ 1 }{ 2 }yx ^{-1}+2y\]

OpenStudy (credmond):

Oh, right! Okay!

OpenStudy (credmond):

So then plug in 1,1?

OpenStudy (anonymous):

\[\frac{ \partial f }{ \partial x } = 4\frac{ \partial }{ \partial x }(x)+\frac{ 1 }{ 2 }\frac{ \partial }{ \partial x }(yx ^{-1})+2\frac{ \partial d }{ \partial dx }(y)\]

OpenStudy (anonymous):

so the partial derivatie w.r.t to x of x is simply 1 right? we are simply breaking the terms

OpenStudy (credmond):

How did you get 1? Sorry..

OpenStudy (anonymous):

if i had y=x what is the derivative dy/dx

OpenStudy (credmond):

1?

OpenStudy (anonymous):

yeah you are just using that concept for a function with more than one variable

OpenStudy (anonymous):

\[\frac{ \partial f }{ \partial x } = \frac{ \partial }{ \partial x }(4x-\frac{ 2y }{ 4x }+2y)=4+\frac{ y }{ 2x ^{2} }\]

OpenStudy (anonymous):

\[\frac{ \partial f }{ \partial y } = -\frac{ 1 }{ 2x }+2\]

OpenStudy (credmond):

Oh okay! Great! Thank you so much!!

OpenStudy (anonymous):

simply plug in (1,1) in each

OpenStudy (credmond):

Right! Thank you!